OpenStudy (anonymous):

This question is asking me to find the maximum y-value on the graph of y=f(x). f(x)=-x^2 + 2x+5 The maximum y-value is y=?

OpenStudy (jennychan12):

are you in calculus?

OpenStudy (anonymous):

Please help me with this problem. The eqution is to help me find the maximum y-value on the graph of y=f(x). f(x)= -x^2 + 2x+5. The maximum y-value is y=? Please help me

OpenStudy (jennychan12):

sorry i just wanna know if you're in calculus cuz if u are, there's an easier way to find the maximum.

OpenStudy (anonymous):

No algerbra 1. This is my last class to get my associate degree. I am so scared. I am not good at math. How are you?

OpenStudy (jennychan12):

it's ok. f(x) = ax^2+bx+c f(x)= -x^2 + 2x+5 so.. a=-1, b=2 and c=5 use x= -b/2a after you find what x equals, plug in into the equation to find the y (f(x)) value.

OpenStudy (anonymous):

Jenny Chan can you help please

OpenStudy (jennychan12):

i did?

OpenStudy (anonymous):

I got 25. Is that right

OpenStudy (jennychan12):

check your math. i got 6

OpenStudy (anonymous):

I really need help with this. I am no good at math and at this point so confused can anyone help me solve my previous question. I need the maximum y-value on the graph of y=f(x). f(x)=-x^2+2x+5. What is the maxzimum y-value. The answer 6 is coming up wrong

OpenStudy (anonymous):

Jenny it is telling me the answer is incorrect????

OpenStudy (jennychan12):

your answer (25) is wrong. it's supposed to be 6

OpenStudy (anonymous):

I see what I did wrong. I did not put a - in front of the x. Thank you

OpenStudy (jennychan12):

no probs